OpenStudy (anonymous):

the radius of circle O is 20, and OC= 8. find AB.

OpenStudy (anonymous):

state complete question .. or draw/attach.. if there is a diagram

OpenStudy (anonymous):

given OC=8 OA=20 AB=?

OpenStudy (anonymous):

here AB = 2AC also use pythogarass theorem for rt angled triangle OAC

OpenStudy (anonymous):

now OA^2 = OC^2 +AC^2 so 400 = 64 + AC^2

OpenStudy (anonymous):

OA = radius of circle...